Shift summary for owl shift 7/16/99

Robert Tschirhart (tsch@fnal.gov)
Fri, 16 Jul 1999 06:43:10 -0500

Shift Summary for owl shift, 07/16/99

On Shift: Phillippe Gouffon, BobT

Hours of beam available at all intensities: 2h 15min
Hours of beam available at nominal intensity: 2h 15min
Hours of E832 data (special runs not included): 1h
Hours of special runs (laser scans, muon runs...): 45min

Accounting of lost time: Trigger/DAQ problems, Main Injector vacuum leak.

========================================================================

Runs ended this shift:

Run 13917 - BEAM_832 (4.9 M events)
Run 13918 - Laser Scan

Tapes brought to FCC this shift: none
========================================================================

Accesses during shift: none
========================================================================

Overview of shift: Inherited fairly smooth 832 running at 6.8E12 from the
evening. The prospects for the shift rapidly fell
apart when serious Trigger/DAQ hangups started to
occur at 01:00. These problems were largely resolved
by 03:00, only to witness the departure of beam to
a Main Injector vacuum leak that continued through
to the end of the shift.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13917 00:16 Stream #5 missing FFFF word, KFREND fixes.

13917 ~80 01:04 Trigger hung, no explicit busies indicated.
Trigger state error: Err11:IDLE "L2-Busy".
Multiple KFRENDs, repeater power cycling
and standard voodoo can't fix it. RickK
paged, and arrives on the scene shortly.

13917 80-133 02:30 Extensive debugging traces problem to spurious
Level-2 busy signals from the KQ/Banana AZTEC
readout modules, (RR167, NIM6, LRS4616).
These spurious busies confound the
trigger state machine, which consequently
freezes, leading to a hung trigger. Problem
vanishes when X-view Level-2 busy sources are
removed. But wait, aren't the X-view busies
disabled via software since they are not in
the trigger?? Yes indeed, we verify that
the X-view busy input spigots are really
disabled. So one of the two following
conditions must have been true:

a) The X and Y view busies were correctly cabled
and the problem was actually in the Y-view and
it disappeared when cables were manipulated.

b) The X and/or Y view busies were incorrectly
cabled w/r to software, and X busies were
connected to Y inputs. Unfortunately we
didn't have the presence of mind to actually
record how it was cabled in detail before the
busy sources were removed.

In any case, the X-view Level-2 busy sources are
now removed, and the Y-view is mapped correctly
w/r to software and all is well.

13917 02:36 No beam, Main Injector vacuum leak at MI-40,
no estimate. We end the run and setup for
a laser scan.

13918 03:00 Beam stop in, Laser Scan. Three passes completed,
run ends successfully.

01:00 - Off and on, Phillippe labeled and preped about
05:00 120 tapes.

05:30 Gas check list completed, Blue Ice blocks replaced
around refrig.

========================================================================
Follow up Notes on Previous Shift

Keep an eye on CsI 3114, 961, and Ring Counter trigger timing.
None of this was completed due to DAQ/accelerator problems. .

========================================================================
Items to Follow up on Next Shift

Keep an eye on CsI 3114, 961, and Ring Counter timing. Also watch
the temperture of the alcohol and moderate with blue ice as necessary.
Watch out for Err11:IDLE "L2-Busy" trigger hung states, which might
be the continuing saga of Level-2 busies.